Circular No. 6299

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 B2 T. Nakamura, National Astronomical Observatory, Tokyo, Japan; and

S. Nakano, Sumoto, report the visual discovery of another comet by Yuji Hyakutake. The following observations are available:

1996 UT  α_{2000}  δ_{2000}  m_{1} Observer
Jan. 30.826  14^{h}31.^{m}2  −24^{o}55′  11.0 Hyakutake
31.74632 14 32 03.^{s}65 −24 56 00′′.7  Kojima
31.74872 14 32 03.87 −24 56 00.9 11.9 Ikari
31.75363 14 32 03.94 −24 56 01.0  Kojima
31.75710 14 32 04.18 −24 56 00.4 11.7  _{”}
31.75712 14 32 04.24 −24 56 01.2 12.0 Ikari

Y. Hyakutake (Hayato-machi, Aira-gun, Kagoshima-ken). 25×150 binocu-

lars. Comet diffuse with central condensation; coma diameter ∼ 2′.5. T. Kojima (YGCO Chiyoda Observatory). 0.25-m f/6.0 reflector + CCD.

Coma diameter ∼ 2′.5, very strong condensation and short tail extended toward west-northwest. Communicated by Nakano.

Y. Ikari (Ohtsu). 0.25-m reflector + CCD. Communicated by Nakano.

R. H. McNaught, Siding Spring, reports the comet as moderately condensed with a 4′ coma, m_{1} = 10 (20×120 binoculars) on Jan. 31.747 UT.

COMET C/1996 B1 (SZCZEPANSKI) Further total magnitude estimates: Jan. 27.823 UT, 11.5 (T. Osawa,

Kawachi-cho, Tochigi-ken, Japan, 0.12-m f/6 reflector; independent photographic discovery); 31.16, 9.3 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector, visual; coma diameter 5′).

COMET 67P/CHURYUMOV-GERASIMENKO Total magnitude estimates (cf. IAUC 6289): Jan. 14.75 UT, 12.0 (H.

Mikuž, Miren, Slovenia, 0.20-m reflector); 17.15, 11.2 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 23.13, 11.4 (Hale).

COMET 65P/GUNN Total visual magnitude estimates by Hale (0.41-m reflector): Jan. 21.54

UT, 13.7; 24.53, 13.7; 25.51, 13.7.

1996 January 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green